solve algebraically
2x^3-x^2-18x+9=0

Answer by stanbon(75887)

solve algebraically
2x^3-x^2-18x+9=0
--------------------------------
I graphed it to find a zero at x=-3.
----------------
Using synthetic division find the other zeroes:
-3)....2....-1....-18....9
.........2....-7.....3...|..0
Remainder = 0 because x=-3 is a zero of the polynomial
Quotient: 2x^2-7x+3
-------------------------
Factor 2x^2-7x+3 using the AC Method:
2x^2-6x-x+3 =0
2x(x-3)-(x-3)=0
(x-3)(2x-1)=0
x =3 or x=1/2
------------------
The zeroes of the orignal equation are
x=-3, x=3, x=1/2
==========================
Cheers,
Stan H.